I have looked everywhere in trying to solve these problems! How would you solve these problems step by step?¿
12-n/5 = 21-n-29
AND
21 = 1 -(4a+2)

Answer by checkley71(8403)

12-N/5=21-N-29 05 12-N/5=-N-8 OR (60-N)/5=-N-8 OR 60-N=-5N-40 OR
100=-4N OR 100/-4=N OR -25=N
PROOF 12-(-25)/5=21-(-25)-29 OR 12+25/5=21+25-29 OR 12+5=46-29 OR 17=17.
21=1-(4A+2) OR 21=1-4A-2 OR 21-1+2=-4A OR 22=-4A OR 22/-4=A OR -5.5=A
PROOF 21=1-((4(-5.5))+2) OR 21=1-((-22)+2) OR 21=1-(-22+2) OR 21=1-(-20) OR
21=1+20 OR 21=21.